Please do FULL problem
A centrifuge in a medical laboratory rotates at a
Answer #1

wi = 2600 rev/min = 3600*2pi/60 rad/s = 376.8 rad/s

wf = 0

θ = 50 turns = 50*2pi = 100pi = 314 radians

From 3rd equation of motion

wf^2 = wi^2 + 2αθ

0 = 376.8^2 + 2α*314

=> α = -226.08 rad/s^2

Hence the magntiude of angular acceleration is 226.08 rad/s^2